# Generic torsion-free groups and Rubin actions

### v1

@@ 3. Some combinatorial group theory
In a Van Kampen diagram $\Delta$ over $H$, we call cells in the interior of $\Delta$ tiles. Tiles coming from relations in $\mathcal{R}_{i}$ will be called $i$–tiles, for $i\in\{1,2\}$. A maximal connected (without cut vertices) union $\Xi$ of $i$–tiles for a fixed value of $i$ will be called an $i$–region. Observe that if a $1$–tile shares an edge with a $2$–tile then the label of that edge is $g$.
If $\Delta$ is a disk diagram over $H$ (i.e. a Van Kampen diagram with no cut edges and no cut vertices) and if $\Xi$ is an $i$–region, then we will say that $\Xi$ is separating if the complement of $\Xi$ in $\Delta$ contains at least two components $\Theta_{1}$ and $\Theta_{2}$ which meet $\partial\Delta$; even if $\Xi$ is nonseparating, $\Xi$ itself may fail to be simply connected and may still topologically separate $\Delta$. Since $\Delta$ is homeomorphic to a disk, an easy combinatorial topology argument shows that $\Delta$ admits at least one non-separating $i$–region, for some $i\in\{1,2\}$. If $\Xi$ is a non-separating $i$–region in $\Delta$ then $\Xi$ meets $\partial\Delta$ in an arc.
@@
Assume the contrary, and let $w(g,h_{2}^{-1}h_{1})$ be a word that expresses a nonzero power $(h_{2}^{-1}h_{1})^{n}$ as a product of conjugates of $g$ and its inverse by elements of $K$. Observe that the expression $(h_{2}^{-1}h_{1})^{n}=w(g,h_{2}^{-1}h_{1})$ is equivalent to some other word $w^{\prime}(g,h_{2}^{-1}h_{1})$ representing the identity in $H$ and having exponent sum $n$ in $h_{2}^{-1}h_{1}$. We therefore let $\Delta$ be a reduced minimal area Van Kampen diagram for a word $w(g,h_{2}^{-1}h_{1})$ which has nonzero exponent sum in $h_{2}^{-1}h_{1}$, and which has a minimal number of occurrences of $h_{2}^{-1}h_{1}$. Generally, $\Delta$ may not be a disk diagram and so may have separating vertices and edges, though there will be at least one positive area disk diagram $\Delta_{0}\subseteq\Delta$.
Observe first that $\partial\Delta_{0}$ must have at least one occurrence of $h_{2}^{-1}h_{1}$, since otherwise $\Delta_{0}$ proves that $g$ has finite order in $H$, which is not the case. Choose a non-separating $i$–region $\Xi$, which we may assume without loss of generality is a $1$–region. Observe that the label of the arc $\partial\Xi\cap\partial\Delta_{0}$ must be of the form
@@
for some nonzero value of $n$. Indeed, otherwise the label of the arc $\partial\Xi\cap\partial\Delta_{0}$ must be $g^{n}$; since the remainder of $\partial\Xi$ coincides with arcs in the boundaries of $2$–regions, we conclude that $\partial\Xi$ is a power of $g$, which is not the case since $G$ is torsion-free and because we assumed the set $\mathcal{R}$ to consist of reduced and cyclically reduced words.
@@ 4. Forcing algebraic disjointness through Banach–Mazur games
Recall that a generic property in this space is a property for which there is a comeager subspace in which all groups satisfy the property. The following is direct consequence of Theorems 1.2.1. and 1.1.2 in  [2]; since a countable conjunction of generic properties is a generic property, in this section we shall assume that a generic torsion-free group satisfies all the properties listed in this theorem:
@@
There is a comeager set $\mathcal{X}$ in the space of enumerated torsion free groups such that for all $G\in\mathcal{X}$:
@@
    All nontrivial elements $1\neq g\in G$ are conjugate. In particular, $G$ is simple.
@@
Let $G$ be a generic countable torsion–free group. Then for all pairs of nontrivial elements $f,g\in G$ such that $[f,g]=1$, we have that $f$ is algebraically disjoint from $g$. In particular, $S_{f}=C_{G}(f)$ for each $f\in G$.
The conclusion of Proposition 4.2 also holds for generic countable groups, which is to say without the requirement that $G$ be torsion-free.
@@
We play a Banach–Mazur game on the space of countable torsion-free groups. In order to prove that a generic group $G$ has the desired properties, it suffices to prove that the first player (player A) has a winning strategy which forces the compiled group to satisfy that for all pairs of nontrivial elements $f,g\in G$ such that $[f,g]=1$, we have that $f$ is algebraically disjoint from $g$. Since the space of countable torsion-free groups is closed and therefore Polish, it suffices to show that at any stage of the game, the conditions played so far are compatible with torsion-freeness.
@@ 1. Case 1
    : It has been forced by now that $[f_{n},g_{n}]=1$. We consider the following subcases:
@@
      : It has been forced by now that $[f_{n},h_{n}]=1$. In this case, player A plays an empty system as a move.
@@
      : It has been forced by now that $[f_{n},h_{n}]\neq 1$. Note that in this case, $h_{n}$ cannot be a power of $g_{n}$ in $\Lambda$. Let $a,b$ be numbers that have not appeared in a move so far. In this case, player $A$ plays the following system:
      $[g_{n},a]=1\qquad[g_{n},b]=1$
      $[a,[b,h_{n}]]\neq 1\qquad[g_{n},[a,[b,h_{n}]]]=1$
@@
      : If neither $[f_{n},h_{n}]=1$ nor $[f_{n},h_{n}]\neq 1$ has been forced so far, player A plays $[f_{n},h_{n}]=1$.
@@ 2. Case 2
    : It has been forced by now that $[f_{n},g_{n}]\neq 1$. In this case, player A plays an empty system as a move.
@@ 3. Case 3
    : If neither $[f_{n},g_{n}]\neq 1$ nor $[f_{n},g_{n}]=1$ has been forced so far, player A plays $[f_{n},g_{n}]\neq 1$.
That the compiled group will have the desired properties is now immediate. ∎
@@
For all nontrivial $g\in G$, we have $C_{G}(S_{g})=\langle g\rangle$. In particular, for each $g_{1},\ldots,g_{n}\in G$ with $n\geq 1$, it holds that
@@
Divisibility is guaranteed by an application of Theorem 4.1. Now, let $g\in G$ be given, and let $1\neq k\in G$ be arbitrary element which is not a power of $g$. Then, player A builds an element $h$ that commutes with $g$ but no nontrivial power of $h$ commutes with $k$ (by building a suitable free product with amalgamation, for instance). This can be done at each stage of the game with all elements which have been played up to that point. Then, by Proposition 4.2 we will have $h^{12}\in S_{g}$ and so $k\notin C_{G}(S_{g})$. Since $g\in C_{G}(S_{g})$ by definition, we see that $\langle g\rangle$ coincides with $C_{G}(S_{g})$. The subsequent claim then follows from the fact that a finite intersection of infinite cyclic groups is trivial or infinite cyclic. ∎
@@ 5. Proof of the main result
Proof.
Suppose that $G$ is a generic torsion-free countable group. By Proposition  4.2, we may assume that all nontrivial elements of $G$ are conjugate, and that if $f,g\in G\setminus\{1\}$ commute then $f$ is algebraically disjoint from $g$. Suppose that
@@
is a homomorphism whose image $\Gamma$ is locally moving. Since $G$ is simple, we have that $G\cong\Gamma$, and so algebraic disjointness in $G$ propagates to algebraic disjointness in $\Gamma$. Since each nontrivial element $f$ is algebraically disjoint from itself, by Proposition 2.1, the elements $f^{12}$ and $f^{12}$ would have disjoint supports as elements of $\Gamma$, which is impossible.
@@
To see that the Rubin poset of the group $G$ consists of cyclic subgroups of $G$ ordered by inclusion with bi-infinite chains and infinite antichains, we simply quote Corollary 4.4 and Corollary 4.3. Proposition 4.2 shows that two nontrivial elements of $G$ commute if and only if they are algebraically disjoint in $G$. ∎

### v2

@@ 3. Some combinatorial group theory
In a Van Kampen diagram $\Delta$ over $H$, we call cells in the interior of $\Delta$ tiles. Tiles coming from relations in $\mathcal{R}_{i}$ will be called ii–tiles, for $i\in\{1,2\}$. A maximal connected (without cut vertices) union $\Xi$ of ii–tiles for a fixed value of $i$ will be called an ii–region. Observe that if a 11–tile shares an edge with a 22–tile then the label of that edge is $g$.
If $\Delta$ is a disk diagram over $H$ (i.e. a Van Kampen diagram with no cut edges and no cut vertices) and if $\Xi$ is an ii–region, then we will say that $\Xi$ is separating if the complement of $\Xi$ in $\Delta$ contains at least two components $\Theta_{1}$ and $\Theta_{2}$ which meet $\partial\Delta$; even if $\Xi$ is non separating, $\Xi$ itself may fail to be simply connected and may still topologically separate $\Delta$.
Since $\Delta$ is homeomorphic to a disk, an easy combinatorial topology argument shows that $\Delta$ admits at least one non-separating ii–region, for some $i\in\{1,2\}$. Indeed, let $\Delta_{0}$ be an $i$-region in $\Delta$. Note that $\Delta_{0}$ must share an edge with $\partial\Delta$, since any edge shared by a 11–region and a 22–region is labeled by $g$; thus, if $\Delta_{0}$ were contained entirely in the interior of $\Delta$ then the boundary label of $\Delta_{0}$ would be a power of $g$. This power of $g$ cannot be nonzero since $G$ is torsion-free and non-trivial. If the power of $g$ were zero then we could cut $\Delta_{0}$ out of $\Delta$ and identify edges in $\partial\Delta_{0}$ to get a smaller area (i.e. fewer tiles) disk diagram $\Delta^{\prime}$ with the same boundary label as $\Delta$.
Writing $\Delta$ as a union $\mathcal{S}$ of distinct 11–regions and $\mathcal{T}$ of distinct 22–regions, we thus conclude that each $\Omega\in\mathcal{S}$ either separates $\Delta$ into two components, or $\Omega$ is non-separating. Every element of $\mathcal{S}\setminus\Omega$ and every element of $\mathcal{T}$ must lie in a component of $\Delta\setminus\Omega$; a similar argument holds for $\Omega\in\mathcal{T}$. By an easy induction, we conclude that there is a non-separating ii–region for some $i$.
Note that if $\Xi$ is a non-separating ii–region in $\Delta$ then $\Xi$ meets $\partial\Delta$ in an arc. With these observations, we are ready to proceed.
@@
Assume the contrary, and let $w^{\prime}(g,h_{2}^{-1}h_{1})$ be a word that expresses a nonzero power $(h_{2}^{-1}h_{1})^{n}$ as a product of conjugates of $g$ and its inverse by elements of $K$. Observe that the expression $(h_{2}^{-1}h_{1})^{n}=w^{\prime}(g,h_{2}^{-1}h_{1})$ is equivalent to some other word $w(g,h_{2}^{-1}h_{1})$ representing the identity in $H$ and having exponent sum $n$ in $h_{2}^{-1}h_{1}$. We therefore let $\Delta$ be a reduced minimal area Van Kampen diagram for a word $w(g,h_{2}^{-1}h_{1})$ which has nonzero exponent sum in $h_{2}^{-1}h_{1}$, and which has a minimal number of occurrences of $h_{2}^{-1}h_{1}$. Generally, $\Delta$ may not be a disk diagram and so may have separating vertices and edges, though there will be at least one positive area disk diagram $\Delta_{0}\subseteq\Delta$.
Observe first that $\partial\Delta_{0}$ must have at least one occurrence of $h_{2}^{-1}h_{1}$, since otherwise $\Delta_{0}$ proves that $g$ has finite order in $H$, which is not the case. Choose a non-separating ii–region $\Xi$, which we may assume without loss of generality is a 11–region. Observe that the label of the arc $\partial\Xi\cap\partial\Delta_{0}$ must be of the form
@@
for some nonzero value of $n$. Indeed, otherwise the label of the arc $\partial\Xi\cap\partial\Delta_{0}$ must be $g^{n}$; since the remainder of $\partial\Xi$ coincides with arcs in the boundaries of 22–regions, we conclude that $\partial\Xi$ is a power of $g$, which is not the case since $G$ is torsion-free and because we assumed the set $\mathcal{R}$ to consist of reduced and cyclically reduced words.
@@ 4. Forcing algebraic disjointness through Banach–Mazur games
Recall that a generic property in this space is a property for which there is a comeager subspace in which all groups satisfy the property. The following is direct consequence of Theorems 1.2.1. and 1.1.2 in  [2]; since a countable conjunction of generic properties is again a generic property, in this section we shall assume that a generic torsion-free group satisfies all the properties listed in this theorem:
@@
There is a comeager set $\mathcal{X}$ in the space of enumerated torsion-free groups such that for all $G\in\mathcal{X}$:
@@
    All non-trivial elements $1\neq g\in G$ are conjugate. In particular, $G$ is simple.
@@
Let $G$ be a generic countable torsion–free group. Then for all pairs of non-trivial elements $f,g\in G$ such that $[f,g]=1$, we have that $f$ is algebraically disjoint from $g$. In particular, $S_{f}=C_{G}(f)$ for each $f\in G$.
@@
We play a Banach–Mazur game on the space of countable torsion-free groups. In order to prove that a generic group $G$ has the desired properties, it suffices to prove that the first player (player A) has a winning strategy which forces all the compiled groups to satisfy that for all pairs of non-trivial elements $f,g\in G$ such that $[f,g]=1$, we have that $f$ is algebraically disjoint from $g$. Since the space of countable torsion-free groups is closed and therefore Polish, it suffices to show that at any stage of the game, the conditions played so far are compatible with torsion-freeness.
@@ 1. Case 1
    : It has been forced by now that $f_{n},g_{n}]=1$. We consider the following subcases:
@@
      : It has been forced by now that $f_{n},h_{n}]=1$. In this case, player A plays an empty system as a move.
@@
      : It has been forced by now that $f_{n},h_{n}]\neq 1$. Note that in this case, $h_{n}$ cannot be a power of $g_{n}$ in $\Lambda$. Let $a,b$ be numbers that have not appeared in a move so far. In this case, player $A$ plays the following system:
      $g_{n},a]=1\qquad[g_{n},b]=1$
      $a,[b,h_{n}]]\neq 1\qquad[g_{n},[a,[b,h_{n}]]]=1$
@@
      : If neither $f_{n},h_{n}]=1$ nor $f_{n},h_{n}]\neq 1$ has been forced so far, player A plays $f_{n},h_{n}]=1$.
@@ 2. Case 2
    : It has been forced by now that $f_{n},g_{n}]\neq 1$. In this case, player A plays an empty system as a move.
@@ 3. Case 3
    : If neither $f_{n},g_{n}]\neq 1$ nor $f_{n},g_{n}]=1$ has been forced so far, player A plays $f_{n},g_{n}]\neq 1$.
That all the compiled groups will have the desired properties is now immediate. ∎
@@
For all non-trivial $g\in G$, we have $C_{G}(S_{g})=\langle g\rangle$. In particular, for each $g_{1},\ldots,g_{n}\in G$ with $n\geq 1$, it holds that
@@
Divisibility is guaranteed by an application of Theorem 4.1. Now, let $g\in G$ be given, and let $1\neq k\in G$ be arbitrary element which is not a power of $g$. Then, player A builds an element $h$ that commutes with $g$ but no non-trivial power of $h$ commutes with $k$ (by building a suitable free product with amalgamation, for instance). This can be done at each stage of the game with all elements which have been played up to that point. Then, by Proposition 4.2 we will have $h^{12}\in S_{g}$ and so $k\notin C_{G}(S_{g})$. Since $g\in C_{G}(S_{g})$ by definition, we see that $\langle g\rangle$ coincides with $C_{G}(S_{g})$. The subsequent claim then follows from the fact that a finite intersection of infinite cyclic groups is trivial or infinite cyclic. ∎
@@ 5. Proof of the main result
Proof of Theorem 1.2.
Suppose that $G$ is a generic torsion-free countable group. By Proposition  4.2, we may assume that all non-trivial elements of $G$ are conjugate, and that if $f,g\in G\setminus\{1\}$ commute then $f$ is algebraically disjoint from $g$. Suppose that
@@
is a homomorphism whose image $\Gamma$ is locally moving. Since $G$ is simple, we have that $G\cong\Gamma$, and so algebraic disjointness in $G$ propagates to algebraic disjointness in $\Gamma$. Since each non-trivial element $f$ is algebraically disjoint from itself, by Proposition 2.1, the elements $f^{12}$ and $f^{12}$ would have disjoint supports as elements of $\Gamma$, which is impossible.
@@
To see that the Rubin poset of the group $G$ consists of cyclic subgroups of $G$ ordered by inclusion with bi-infinite chains and infinite antichains, we simply quote Corollary 4.4 and Corollary 4.3. Proposition 4.2 shows that two non-trivial elements of $G$ commute if and only if they are algebraically disjoint in $G$. ∎